# Non-explosive pre-supernova feedback in the COLIBRE model of galaxy formation

### v1

@@ 6 Discussion
As discussed previously, the primary motivation for our NEPS feedback module is to model early physical processes that begin to regulate star formation before the onset of SN feedback, and thus improve numerical convergence of simulations of galaxy formation, particularly at high resolution. As demonstrated in our feedback-free experiments (Figs. 78, and 9), simulations that resolve the cold, dense phase of the ISM are prone to runaway gravitational fragmentation and resolution-dependent star formation histories. Our results show that the implementation of a physically motivated non-explosive pre-supernova feedback module in the COLIBRE model provides an effective solution to this problem, enabling more robust and self-regulated galaxy growth over time. Here we discuss the physical implications of our experiments, the limitations of our model, and future steps.
@@ Acknowledgements
ABL acknowledges support by the Italian Ministry for Universities (MUR) program “Dipartimenti di Eccellenza 2023-2027” within the Centro Bicocca di Cosmologia Quantitativa (BiCoQ), and support by UNIMIB’s Fondo Di Ateneo Quota Competitiva (project 2024-ATEQC-0050). SP acknowledges support by the Austrian Science Fund (FWF) through grant 10.55776/V982. EC and FH acknowledge funding from the Netherlands Organization for Scientific Research (NWO) through research programme Athena 184.034.002. JT acknowledges support of a STFC Early Stage Research and Development grant (ST/X004651/1).
@@ Appendix A Stellar winds kick velocity
Fig. 20 explores the impact of the strength of the stochastic velocity kick, $\Delta v_{0}$, on the star formation rate (SFR) for an intermediate-mass disk ($M_{\rm d}=10^{10}\rm\ M_{\odot}$) at m5 resolution, including only momentum from stellar winds. The results demonstrate two key aspects of our model.
@@
First, the SFR is only weakly dependent on the chosen kick velocity; simulations spanning nearly two orders of magnitude in $\Delta v_{0}$ produce comparable star formation histories. Second, a subtle, non-monotonic trend is visible. Increasing the kick velocity up to 25 km s-1 slightly decreases the average SFR, as expected. However, for values exceeding 25 km s-1, the SFR begins to rise again and become comparable to runs without stellar winds (shown by the grey dashed line). This occurs because the probability of kicking any given gas particle is inversely proportional to $\Delta v_{0}$; for extremely high kick velocities, the momentum is imparted to so few particles that its overall regulatory effect on the disk diminishes. This provides a soft justification for our adoption of $\Delta v_{0}=50$ km s-1 in the fiducial COLIBRE model.
We note, however, that since stellar winds are a sub-dominant component of the full NEPS feedback model, the precise choice of $\Delta v_{0}$ has negligible impact on the final simulation outcomes.
Figure 20: Star formation rate as a function of time for an intermediate-mass ($M_{\rm d}=10^{10}\rm\ M_{\odot}$) gas disk simulated at m5 resolution. These simulations include only momentum feedback from stellar winds. Each line corresponds to a different value of the desired velocity kick, $\Delta v_{0}$, as indicated in the legend. The grey dashed line corresponds to a simulation without stellar winds. The figure shows that the star formation history is only weakly sensitive to the chosen kick velocity.

### v2

@@ 6 Discussion
As discussed previously, the primary motivation for our NEPS feedback module is to model early physical processes that begin to regulate star formation before the onset of SN feedback, and thus improve numerical convergence of simulations of galaxy formation, particularly at high resolution see Table 1 for a summary of convergence for various physical quantities across the simulated models). As demonstrated in our feedback-free experiments (Figs. 78, and 9), simulations that resolve the cold, dense phase of the ISM are prone to runaway gravitational fragmentation and resolution-dependent star formation histories. Our results show that the implementation of a physically motivated non-explosive pre-supernova feedback module in the COLIBRE model provides an effective solution to this problem, enabling more robust and self-regulated galaxy growth over time.
Table 1: Summary of convergence for key physical quantities across the simulated models.
Physical Quantity Model Configuration Status Physical/Numerical Explanation
Mid-plane Gas Density All Models Not Converged The vertical scale height of the cold gaseous disk remains unresolved even at our highest resolution (m4). The maximum density is driven by numerics rather than physics.
Star Formation Histories Feedback-Free Poor Simulations undergo runaway gravitational collapse. Higher resolution resolves smaller, denser clumps, leading to notably higher SFRs.
NEPS (Early Feedback) Good Thermal pressure from H ii regions regulates the collapse of dense gas. The SFR becomes largely independent of numerical resolution once regulated.
Full Model (NEPS + SN) Excellent The synergy between early regulation and SN feedback prevents both initial starbursts and late-time high SFRs, leading to a highly stable, converged SFH.
Gas Temperature Structure Feedback-Free Poor The mass fraction of cold gas (<100<100 K) depends strongly on resolution, driven by resolution-dependent runaway collapse.
NEPS / Full Model Good The inclusion of photoheating establishes a stable three-phase ISM (Cold, Warm, Ionized). The PDF of gas temperatures becomes robust to changes in resolution.
Disk Morphology Feedback-Free Poor Low-resolution runs are artificially stabilized by numerical pressure, while high-resolution runs fragment into small dense clumps.
Full Model (NEPS + SN) Good NEPS feedback prevents the formation of overly dense clusters, leading to distributed SNe that maintain a smooth, homogeneous disk structure across resolutions.
@@
Third, our estimate of the optical depth neglects the detailed subgrid density structure of the gas. The presence of unresolved dense clumps can enhance the coupling of radiation pressure, or conversely, create low-density channels that allow radiation to leak. However, given the pronounced dominance of the thermal H ii region feedback channel we expect that even a moderate increase in the radiation pressure coupling efficiency due to subgrid clumping would not overturn the established hierarchy of feedback mechanisms shown in Fig 15.
@@ Acknowledgements
We acknowledge the anonymous referee for comments and suggestions that helped improve the presentation of our results and further justify our numerical choices. ABL acknowledges support by the Italian Ministry for Universities (MUR) program “Dipartimenti di Eccellenza 2023-2027” within the Centro Bicocca di Cosmologia Quantitativa (BiCoQ), and support by UNIMIB’s Fondo Di Ateneo Quota Competitiva (project 2024-ATEQC-0050). SP acknowledges support by the Austrian Science Fund (FWF) through grant 10.55776/V982. EC and FH acknowledge funding from the Netherlands Organization for Scientific Research (NWO) through research programme Athena 184.034.002. JT acknowledges support of a STFC Early Stage Research and Development grant (ST/X004651/1).
@@ Appendix A Stellar winds kick velocity
As detailed in Sec. 2, our stochastic momentum injection scheme requires defining a target kick velocity, $\Delta v_{0}$, to determine the probability of kicking neighboring particles. While the total momentum delivered to the ISM is fixed by the BPASS model, the choice of $\Delta v_{0}$ determines how this momentum is sampled. A higher velocity implies kicking fewer particles (i.e., lower mass loading) to conserve the total budget. Here, we explore the sensitivity of our results to this parameter.
Figure 20 shows the star formation history of the $10^{10}\,\text{M}_{\odot}$ disk for $\Delta v_{0}$ ranging from $\approx 3$ to $100$ km s-1. The top panel shows results at our lower (m5) resolution. In this regime, the disk is artificially stabilized as discussed in Sec. 4.2. Consequently, the momentum from stellar winds is expected to have a negligible impact, which is indeed what this panel shows; the SFR is only weakly dependent on the chosen kick velocity. Simulations spanning nearly two orders of magnitude in $\Delta v_{0}$ produce comparable star formation histories.
@@
However, a very subtle, non-monotonic trend is visible. Increasing the kick velocity up to 25 km s-1 slightly decreases the average SFR at later times, as expected. Yet, for values exceeding 25 km s-1, the SFR begins to rise again, becoming comparable to runs without stellar winds (shown by the grey dashed line). This occurs because the probability of kicking any given gas particle is inversely proportional to $\Delta v_{0}$; for extremely high kick velocities, the momentum is imparted to so few particles that its overall regulatory effect on the disk diminishes.
The bottom panel of Figure 20 shows a similar test at the higher (m4) resolution, where the disk is numerically unstable and early feedback is essential for regulation. For all values of $\Delta v_{0}$ explored, we confirm that stellar winds have a significant impact on the SFR, suppressing it by a factor of $\approx 2$ at later times. Despite this, a clear trend with $\Delta v_{0}$ emerges. Similar to the trend observed for the m5 counterpart, velocity kicks between $\approx 12$ to $25$ km s-1 maximize the suppression of the SFR. For kick velocities $\lesssim 12$ km s-1 (blue/cyan lines), the winds become increasingly ineffective. This is likely because the velocity imparted to the star-forming clouds is comparable to or less than the local effective sound speed, rendering the kicks dynamically less relevant.
For higher velocities, $\gtrsim 50$ km s-1, the star formation rate increases with the kick velocity. As with the lower resolution case, this occurs because conserving the total momentum with a higher target kick velocity requires a decrease in mass loading. At very high velocities, the scheme kicks only a few particles very hard; these high-speed ejecta may escape the local dense region without efficiently coupling their momentum to the bulk of the cloud.
Despite this second-order non-monotonic behaviour, the variations in SFR between models with different kick velocities are relatively modest. Given that thermal pressure from H ii regions remains the dominant stabilizing mechanism, our fiducial choice of $\Delta v_{0}=50$ km s-1 places our model in the robust, effective regime.
Figure 20: Star formation rate as a function of time for the intermediate-mass disk ($M_{\rm d}=10^{10}\ \text{M}_{\odot}$), varying the stellar winds kick velocity parameter, $\Delta v_{0}$, as indicated in the legend. The total momentum is kept constant in all runs, meaning higher velocities result in fewer particles kicked. The dashed grey line indicates the baseline run without stellar winds. The top panel shows results for the intermediate (m5) resolution, for which the SFR is insensitive to the kick velocity because the disk structure is numerically stabilized. In the bottom panel, we show the high-resolution (m4) counterpart. Here the disk is both physically and numerically unstable. The SFR is maximally suppressed for wind speeds of $\approx 25$ km s-1. Above and below this threshold, the SFR is less supressed, but the differences are small.
Appendix B H ii regions lifetime
[…]